Michelle jogs 12 miles in one hour and 30 mins. Find speed in MPH.



What if she jog 18 miles
: Michelle jogs 12 miles in one hour and 30 mins. Find speed in MPH.



What if she jog 18 miles

Answer by stanbon(19722)

Michelle jogs 12 miles in one hour and 30 mins. Find speed in MPH.
Rate = distance/time
Rate = 12 mi/ (3/2) hr = 24/3 = 8 mph
============================================
What if she jogs 18 miles?
Rate = distance /time
Rate = 18 mi/ (3/2) hr = 36/3 = 12 mph
=============================================
Cheers,
Stan H.